# ZTF-EP-25-07

### v1

@@ Abstract
The search for regular black holes with nonlinear electromagnetic fields has sprouted numerous candidates, each exhibiting certain virtues but often accompanied by significant drawbacks. We demonstrate that Komar mass, electric charge and magnetic charge are mutually dependent in regular black holes with nonlinear electromagnetic fields, defined by Lagrangian which is a function of both electromagnetic invariants, $F_{ab}F^{ab}$ and $F_{ab}{\star F}^{ab}$, regardless of the specific weak field limit of the theory. Also, we generalize one of the key no-go theorems by showing that static, spherically symmetric, electrically charged black holes in a theory respecting the relaxed Maxwellian weak field limit do not admit a bounded Kretschmann scalar. Finally, we address one of the long-standing niche questions, whether regular black hole solutions can exist when both electric and magnetic charges are present, by constructing an exotic family of regular dyonic black holes with nonlinear electromagnetic fields in theories respecting the Maxwellian weak field limit. Mounting evidence suggests that regularizing black holes through simplistic nonlinear extensions of Maxwell’s electromagnetism entails a high cost in the form of unorthodox theoretical assumptions.
@@ 1 Introduction
An ever-expanding line of research in theoretical gravitational physics originates from the deceptively simple, yet heavily loaded, question “Do regular black holes exist?”. Without any additional disclaimer, the answer is resounding yes, followed by dramatic however. A cheap way to produce innumerable examples is based on a simple trick, sometimes referred to as the “Synge G-method” [EG24]: Insert spacetime metric $g_{ab}$, regular in some sense which you consider appropriate (geodesic completeness, bounded curvature invariants, etc.) into gravitational field equation and proclaim “right-hand side” to be energy-momentum tensor of some physical matter. A seminal work of this approach in the context of the search for regular black holes is Bardeen’s two-page proceedings paper [Bar68], featuring a brief discussion of a modified Schwarzschild metric (a more recent resurrection of the idea has appeared in [SV19]). As was already noted by Bardeen, a basic constraint on the corresponding energy-momentum tensor is given by the various energy conditions [Mae22]. A far more nontrivial challenge lies in constructing a complete Lagrangian (for all the fields and their interaction), such that the corresponding gravitational-matter field equations admit a solution describing a regular black hole spacetime. All such candidates may be sorted based on the extent to which they depart from established physical theories and ultimately put under the scrutiny of experiments.
@@
Over the past quarter of century, a significant effort has been put into investigation of gravitating electromagnetic fields. As we already know that conventional Einstein–Maxwell field equations lead to Kerr–Newmann family of singular black hole solutions, a survey had to be spread over various modified theories. First breakthrough was made by Ayón-Beato and García [ABG98], but it was soon realized that this solution suffers from a serious drawback [Bro00], stemming from multivaluedness of the Lagrangian. A couple of years later, the same authors managed [ABG00] to find an electromagnetic Lagrangian which reproduces Bardeen’s metric as a magnetically charged black hole. Again, this theory exhibits one technical limitation and one physically unrealistic feature, both of which persist in numerous subsequent proposals. First, Ayón-Beato–García electromagnetic Lagrangian has singular behaviour in the weak field limit (in the sense which will be defined more precisely below), making it difficult to connect with the conventional, well-tested Maxwell’s electromagnetism. In fact, a series of no-go theorems [Bro01, BSJ22a] constraint theories which admit regular black hole solutions and satisfy Maxwell’s weak field limit. Yet, it is still possible to find theories, via Lagrangian reverse engineering, with a satisfactory weak field limit [BFJS24], admitting regular, magnetically charged black holes. Second, Ayón-Beato–García electromagnetic Lagrangian does not have independent black hole mass and magnetic charge [^1] This point is somewhat muddled in the original paper [ABG00], with the mass and the magnetic charge appearing among the Lagrangian parameters. As will be shown in the Appendix B, this technical issue may be resolved with appropriate redefinitions, at the cost of making the magnetic charge a function of mass (or vice versa).. As far as we are aware of, this feature is common for all the regular black holes found in theories with minimally coupled nonlinear electromagnetic Lagrangians. One of the objectives of the present work is to clarify how universal is this phenomenon.
@@ 2 Panoramic view of gravitating nonlinear electromagnetic fields
Nonlinear electromagnetic (NLE) theories are, in general, defined by Lagrangians which contain electromagnetic tensor $F_{ab}$ and its Hodge dual ${\star F}_{ab}\mathrel{\mathop{:}}=\epsilon_{abcd}F^{cd}/2$. Basic class of NLE Lagrangians are functions of various contractions of $F_{ab}$ and ${\star F}_{ab}$ (no additional covariant derivatives). It can be shown [EU14] that such NLE Lagrangians are in fact functions $\mathscr{L}(\mathcal{F},\mathcal{G})$ of two electromagnetic invariants,
@@
Historically, NLE theories appeared back in the 1930s, with two prominent examples: the Born–Infeld Lagrangian [Bor34, BI34], constructed with the specific aim to cure the inconsistencies of Maxwell’s electromagnetism associated with the infinite self-energy of the point charges, and the Euler–Heisenberg one-loop QED correction to Maxwell’s Lagrangian [HE36]. The ensuing century brought us plethora of proposed NLE theories [Sor22, Bro22]. Nonlinearities of the electromagnetic field are being tested by the ATLAS Collaboration [A+17, EMY17, NAM19] and new generations of the high-power lasers at the Extreme Light Infrastructure [T+20] (see also [BR13, FBR16]).
@@
Both definitions encapsulate some notion that the electromagnetic Lagrangian is “Maxwell-like” for weak fields. The difference is that for the MWF limit we impose conditions on some neighbourhood of the origin of the $\mathcal{F}$-$\mathcal{G}$ plane (thus, for a broader family of dyonic solutions), whereas for the qMWF limit we only demand that the Lagrangian has an appropriate limit in the absence of magnetic charge, as the electric field is turned off. The MWF limit is a stricter condition than the qMWF limit: while the former implies the latter, the converse implication does not necessarily follow. For example, Born–Infeld and Euler–Heisenberg Lagrangians satisfy MWF limit, while ModMax Lagrangian [Kos20, BLST20] satisfies only qMWF and not the MWF limit. Just to complete the nomenclature, we may denote members of qMWF branch which do not satisfy MWF limit as Lagrangians with the “directional MWF” (dMWF) limit. Any NLE Lagrangian $\mathscr{L}(\mathcal{F},\mathcal{G})$ which fails even at qMWF limit will be referred to as “irregular”. An elementary example of an irregular NLE Lagrangian is the so-called power-Maxwell $\mathscr{L}(\mathcal{F})=C|\mathcal{F}|^{s}$, with constant $C\neq 0$ and the exponent $s\neq 1$.
More complex family of NLE theories are those in which the Lagrangian contains contractions of terms with covariant derivatives, such as $\nabla_{\!a}F_{bc}$, $\nabla_{\!a}{\star F}_{bc}$, $\nabla_{\!a}\nabla_{\!b}F_{cd}$, etc. Members of this branch are collectively referred to as higher-derivative NLE theories, with the closely related Bopp–Landé–Thomas–Podolsky [Bop40, Lan41, LT41, Pod42] and Lee–Wick [LW70] theories serving as the paradigmatic examples (see also [TN14, CdMM+18, Amo25, SP25] and references therein). A natural generalization of the MWF limit for higher-derivative NLE Lagrangians begins with identifying a set of independent invariants $(\mathcal{F},\mathcal{G},\mathcal{I}_{3},\dots,\mathcal{I}_{n})$ that serve as building blocks. One can then say that a Lagrangian of the form $\mathscr{L}(\mathcal{F},\mathcal{G},\mathcal{I}_{3},\dots,\mathcal{I}_{n})$ satisfies an appropriate weak field limit if it is suitably smooth in a neighborhood of the origin in the formal space $\mathds{R}^{n}$ of invariants (we shall not pursue the technical details of this construction here).
@@
In the present work we are mainly focused on the $\mathscr{L}(\mathcal{F},\mathcal{G})$-branch of gNLE theories, with the standard Einstein–Hilbert gravitational action. Using the auxiliary 2-form $\mathbf{Z}\mathrel{\mathop{:}}=-4(\mathscr{L}_{\mathcal{F}}\,\mathbf{F}+\mathscr{L}_{\mathcal{G}}\,{\star\mathbf{F}})$, the corresponding source-free NLE Maxwell’s equations can be written concisely as
@@ 3 Static, spherically symmetric case
In this paper, our main focus is on the static, spherically symmetric, asymptotically flat solutions of the Einstein–NLE field equations. More precisely, spacetime consists of a smooth 4-manifold $M$, with an asymptotically flat Lorentzian metric $g_{ab}$ and the electromagnetic 2-form $F_{ab}$. Furthermore, metric $g_{ab}$ admits a hypersurface orthogonal Killing vector field $k^{a}$, timelike in the asymptotic region, and the group of isometries contains a subgroup isomorphic to $SO(3)$, whose orbits are 2-spheres which foliate hypersurfaces to which $k^{a}$ is orthogonal. On each of the orbit 2-spheres we may define spherical coordinates $(\theta,\varphi)$, such that they are constant along integral curves of a smooth congruence of geodesics orthogonal to the group orbits. Furthermore, the radial coordinate is introduced as the function $r=(A/4\pi)^{1/2}$ of the area $A$ of the orbit 2-spheres. Leaving aside “wormhole” and “horn” geometries (see discussion in [Bro01]), we assume that $\nabla_{a}r\neq 0$ everywhere in the spacetime and radial coordinate takes values on the interval $r\in\left[0,\infty\right$>. Finally, Killing parameter $t$, defined via $k^{a}\nabla_{\!a}t=1$, may be used to complete the coordinate system into canonical $(t,r,\theta,\varphi)$, given that $k^{a}$ and $\nabla^{a}r$ are not collinear. As is well-known, the latter case occurs on Killing horizons (on which $k^{a}$ is lightlike), thus the constructed coordinate system is valid only on domains between adjacent Killing horizons, while for the extension across the horizon one must resort to new coordinates tailored along convenient geodesics (such as Eddington–Finkelstein and Kruskal coordinates). Note that the causal character of vectors $(k^{a},\nabla^{a}r)$ may switch from (timelike, spacelike) in one region to (spacelike, timelike) in other region, so that the names “time coordinate” and “radial coordinate” should be taken with a grain of salt.
@@
Regularity of the metric is defined in a sense that the associated Kretschmann scalar $K\mathrel{\mathop{:}}=R_{abcd}R^{abcd}$ is a bounded continuous real function on the interval $\left[0,\infty\right$>. In particular, we say that a metric has a “regular center” if $\lim_{r\to 0^{+}}K$ is well-defined and finite. Kretschmann scalar $K$ for the metric (7) is a sum of three nonnegative terms,
@@
Note that a bounded Kretschmann scalar implies that left hand sides of both gravitational field equations remain bounded as $r\to 0^{+}$. As was shown by two theorems in [BSJ22b], strictly stationary, regular solutions of the Einstein-NLE field equations, with the energy-momentum tensor respecting the dominant energy condition, have either a trivial or stealth electromagnetic field. Thus, our main focus is on the regular black hole solutions.
@@ 4 Mass and charges entangled
In other words, to ensure that the Kretschmann scalar is bounded, the Komar mass $M$ and the magnetic charge $P$ must be constrained by the equation above. A special, pathological case occurs if $I=0$, implying that we need $M=0$ for regularization. Relation (22) should not come as a surprise, as it was already present in the Ayón-Beato–García’s NLE interpretation of the Bardeen’s black hole (see comments in the Appendix B) and it has appeared in more recent reverse engineered NLE Lagrangians [BFJS24] (see comments in the Appendix A).
@@ 5 Generalized constraints
Let us turn to another kind of constraint on regular black holes with NLE fields. Back in 2001 Bronnikov [Bro01] proved that an NLE Lagrangian $\mathscr{L}(\mathcal{F})$ with the MWF limit does not admit electrically charged, static, spherically symmetric black holes with bounded Ricci scalar $R$ and “Ricci squared” scalar $S\mathrel{\mathop{:}}=R_{ab}R^{ab}$. This result was later generalized [BSJ22a] to a larger class of $\mathscr{L}(\mathcal{F},\mathcal{G})$ Lagrangians. Using the terminology from section 2, in combination with the recent results from [DMS25], it is straightforward to extend and strengthen the conclusion even further, in the form of the following theorem.
@@
This generalization immediately applies to theories respecting qMWF and not necessarily MWF limit, such as ModMax theory. Another perspective may be added here, as ModMax theory belongs to the larger class of NLE theories with vanishing trace of energy-momentum tensor, $T=0$. Einstein’s gravitational field equation with the cosmological constant $\Lambda$ (included here for the sake of generality) implies for any member of such a theory that the Ricci scalar is constant, so that for static, spherically symmetric spacetimes considered here, we have

### v2

@@ Abstract
The search for regular black holes with nonlinear electromagnetic fields has sprouted numerous candidates, each exhibiting certain virtues but often accompanied by significant drawbacks. We demonstrate that Komar mass, electric charge and magnetic charge are mutually dependent in regular black holes with nonlinear electromagnetic fields, defined by a Lagrangian which is a function of both electromagnetic invariants, $F_{ab}F^{ab}$ and $F_{ab}{\star F}^{ab}$, regardless of the specific weak field limit of the theory. Also, we generalize one of the key no-go theorems by showing that static, spherically symmetric, electrically charged black holes in a theory respecting the relaxed Maxwellian weak field limit do not admit a bounded Kretschmann scalar. Finally, we address one of the long-standing niche questions, whether regular black hole solutions can exist when both electric and magnetic charges are present, by constructing an exotic family of regular dyonic black holes with nonlinear electromagnetic fields in theories respecting the Maxwellian weak field limit. Mounting evidence suggests that regularizing black holes through simplistic nonlinear extensions of Maxwell’s electromagnetism entails a high cost in the form of unorthodox theoretical assumptions.
@@ 1 Introduction
An ever-expanding line of research in theoretical gravitational physics originates from the deceptively simple, yet heavily loaded, question “Do regular black holes exist?”. Without any additional disclaimer, the answer is resounding yes, followed by dramatic however. A cheap way to produce innumerable examples is based on a simple trick, sometimes referred to as the “Synge G-method” [EG24]: Insert spacetime metric $g_{ab}$, regular in some sense which you consider appropriate (geodesic completeness, bounded curvature invariants, etc.) into gravitational field equation and proclaim “right-hand side” to be energy-momentum tensor of some physical matter. A seminal work of this approach in the context of the search for regular black holes is Bardeen’s two-page proceedings paper [Bar68], featuring a brief discussion of a modified Schwarzschild metric (a more recent resurrection of the idea has appeared in [SV19]). As was already noted by Bardeen, a basic constraint on the corresponding energy-momentum tensor is given by the various energy conditions [Mae22]. A far more nontrivial challenge lies in constructing a complete Lagrangian (for all the fields and their interaction), such that the corresponding gravitational-matter field equations admit a solution describing a regular black hole spacetime [^1] For example, one approach studies regular black holes with de Sitter cores [Dym92, DK15, Ver25], while more recent proposals consider a signature change at the event horizon [CBDB24, CBDB25] and an infinite tower of higher-curvature Lagrangian corrections in higher-dimensional spacetimes [BCH25].. All such candidates may be sorted based on the extent to which they depart from established physical theories and ultimately put under the scrutiny of experiments.
Another key aspect in assessing the physical viability of regular black holes is understanding their potential dynamical formation. Several scenarios have been proposed in which repulsive effects prevent singular collapse. These include mechanisms motivated by quantum gravity, such as effective corrections from the asymptotic safety program [BMP24] or phenomenological models [ZK10]. Regular black holes may also be produced dynamically through purely classical mechanisms, relying on modified gravitational theories [Pop21, BCHM25] or instabilities in AdS backgrounds, which may trigger the creation of de Sitter patches corresponding to the interiors of the regular black holes [BES22]. Formation mechanisms have also been proposed for a broader class of horizonless compact objects that, based on their observational signatures, could mimic black holes [GRB09, HPR+21, SCHR24]. From a theoretical perspective, these solitoniclike configurations offer a promising avenue for resolving the singularity problem without departing from the framework of general relativity, often using scalar and Proca fields as their matter content. Boson (see [Shn23] for a review) and Proca [BCHR16] stars are thought to arise through collapse accompanied by gravitational cooling [SS94, DGSGHF18], while the formation of gravastars [MM04] appears to rely on finely tuned dust sphere collapse and the expansion of a de Sitter core [JR25].
@@
Over the past quarter of century, a significant effort has been put into investigation of gravitating electromagnetic fields. As we already know that conventional Einstein–Maxwell field equations lead to Kerr–Newmann family of singular black hole solutions, a survey had to be spread over various modified theories. The first breakthrough was made by Ayón-Beato and García [ABG98], but it was soon realized that this solution suffers from a serious drawback [Bro00], stemming from multivaluedness of the Lagrangian. A couple of years later, the same authors managed [ABG00] to find an electromagnetic Lagrangian which reproduces Bardeen’s metric as a magnetically charged black hole. Again, this theory exhibits one technical limitation and one physically unrealistic feature, both of which persist in numerous subsequent proposals. First, the Ayón-Beato–García electromagnetic Lagrangian has singular behaviour in the weak field limit (in the sense which will be defined more precisely below), making it difficult to connect with the conventional, well-tested Maxwell’s electromagnetism. In fact, a series of no-go theorems [Bro01, BSJ22a] constrain theories that admit regular black hole solutions and satisfy Maxwell’s weak field limit. Yet, it is still possible to find theories, via Lagrangian reverse engineering, with a satisfactory weak field limit [BFJS24], admitting regular, magnetically charged black holes. Second, the Ayón-Beato–García electromagnetic Lagrangian does not have independent black hole mass and magnetic charge [^2] This point is somewhat muddled in the original paper [ABG00], with the mass and the magnetic charge appearing among the Lagrangian parameters. As will be shown in the Appendix B, this technical issue may be resolved with appropriate redefinitions, at the cost of making the magnetic charge a function of mass (or vice versa).. As far as we are aware of, this feature is common for all the regular black holes found in theories with minimally coupled nonlinear electromagnetic Lagrangians. One of the objectives of the present work is to clarify how universal this phenomenon is.
@@ 2 Panoramic view of gravitating nonlinear electromagnetic fields
Nonlinear electromagnetic (NLE) theories are, in general, defined by Lagrangians which contain electromagnetic tensor $F_{ab}$ and its Hodge dual ${\star F}_{ab}\mathrel{\mathop{:}}=\epsilon_{abcd}F^{cd}/2$. The basic class of NLE Lagrangians are functions of various contractions of $F_{ab}$ and ${\star F}_{ab}$ (no additional covariant derivatives). It can be shown [EU14] that such NLE Lagrangians are in fact functions $\mathscr{L}(\mathcal{F},\mathcal{G})$ of two electromagnetic invariants,
@@
Historically, NLE theories appeared in the 1930s, with two prominent examples: the Born–Infeld Lagrangian [Bor34, BI34], constructed with the specific aim to cure the inconsistencies of Maxwell’s electromagnetism associated with the infinite self-energy of the point charges, and the Euler–Heisenberg one-loop QED correction to Maxwell’s Lagrangian [HE36]. The ensuing century brought us plethora of proposed NLE theories [Sor22, Bro22, AEM21]. Nonlinearities of the electromagnetic field are being tested by the ATLAS Collaboration [A+17, EMY17, NAM19] and new generations of the high-power lasers at the Extreme Light Infrastructure [T+20] (see also [BR13, FBR16]).
@@
Both definitions encapsulate some notion that the electromagnetic Lagrangian is “Maxwell-like” for weak fields. The difference is that for the MWF limit we impose conditions on some neighbourhood of the origin of the $\mathcal{F}$-$\mathcal{G}$ plane (thus, for a broader family of dyonic solutions), whereas for the qMWF limit we only demand that the Lagrangian has an appropriate limit in the absence of magnetic charge, as the electric field is turned off. The MWF limit is a stricter condition than the qMWF limit: while the former implies the latter, the converse implication does not necessarily follow. For example, the Born–Infeld and Euler–Heisenberg Lagrangians satisfy the MWF limit, while ModMax Lagrangian [Kos20, BLST20] satisfies only qMWF and not the MWF limit. Just to complete the nomenclature, we may denote members of qMWF branch which do not satisfy MWF limit as Lagrangians with the “directional MWF” (dMWF) limit. Any NLE Lagrangian $\mathscr{L}(\mathcal{F},\mathcal{G})$ that fails even at the qMWF limit will be referred to as “irregular”. An elementary example of an irregular NLE Lagrangian is the so-called power-Maxwell $\mathscr{L}(\mathcal{F})=C|\mathcal{F}|^{s}$, with constant $C\neq 0$ and the exponent $s\neq 1$.
A more complex family of NLE theories are those in which the Lagrangian contains contractions of terms with covariant derivatives, such as $\nabla_{\!a}F_{bc}$, $\nabla_{\!a}{\star F}_{bc}$, $\nabla_{\!a}\nabla_{\!b}F_{cd}$, etc. Members of this branch are collectively referred to as higher-derivative NLE theories, with the closely related Bopp–Landé–Thomas–Podolsky [Bop40, Lan41, LT41, Pod42] and Lee–Wick [LW70] theories serving as the paradigmatic examples (see also [TN14, CdMM+18, Amo25, SP25] and references therein). A natural generalization of the MWF limit for higher-derivative NLE Lagrangians begins with identifying a set of independent invariants $(\mathcal{F},\mathcal{G},\mathcal{I}_{3},\dots,\mathcal{I}_{n})$ that serve as building blocks. One can then say that a Lagrangian of the form $\mathscr{L}(\mathcal{F},\mathcal{G},\mathcal{I}_{3},\dots,\mathcal{I}_{n})$ satisfies an appropriate weak field limit if it is suitably smooth in a neighborhood of the origin in the formal space $\mathds{R}^{n}$ of invariants (we shall not pursue the technical details of this construction here).
@@
In the present work, we analyse the $\mathscr{L}(\mathcal{F},\mathcal{G})$-branch of gNLE theories, with the standard Einstein–Hilbert gravitational action. Using the auxiliary 2-form $\mathbf{Z}\mathrel{\mathop{:}}=-4(\mathscr{L}_{\mathcal{F}}\,\mathbf{F}+\mathscr{L}_{\mathcal{G}}\,{\star\mathbf{F}})$, the corresponding source-free NLE Maxwell’s equations can be written concisely as
@@ 3 Static, spherically symmetric case
In this paper, our main focus is on the static, spherically symmetric, asymptotically flat solutions of the Einstein–NLE field equations. More precisely, spacetime consists of a smooth 4-manifold $M$, with an asymptotically flat Lorentzian metric $g_{ab}$ and the electromagnetic 2-form $F_{ab}$. Furthermore, metric $g_{ab}$ admits a hypersurface orthogonal Killing vector field $k^{a}$, timelike in the asymptotic region, and the group of isometries contains a subgroup isomorphic to $SO(3)$, whose orbits are 2-spheres which foliate hypersurfaces to which $k^{a}$ is orthogonal. On each of the orbit 2-spheres we may define spherical coordinates $(\theta,\varphi)$, such that they are constant along integral curves of a smooth congruence of geodesics orthogonal to the group orbits. Furthermore, the radial coordinate is introduced as the function $r=(A/4\pi)^{1/2}$ of the area $A$ of the orbit 2-spheres. Leaving aside “wormhole” and “horn” geometries (see discussion in [Bro01]), we assume that $\nabla_{a}r\neq 0$ everywhere in the spacetime and radial coordinate takes values on the interval $r\in\left[0,\infty\right$>. Finally, Killing parameter $t$, defined via $k^{a}\nabla_{\!a}t=1$, may be used to complete the coordinate system into canonical $(t,r,\theta,\varphi)$, given that $k^{a}$ and $\nabla^{a}r$ are not collinear. As is well known, the latter case occurs on Killing horizons (on which $k^{a}$ is lightlike), thus the constructed coordinate system is valid only on domains between adjacent Killing horizons, while for the extension across the horizon one must resort to new coordinates tailored along convenient geodesics (such as Eddington–Finkelstein and Kruskal coordinates). Note that the causal character of vectors $(k^{a},\nabla^{a}r)$ may switch from (timelike, spacelike) in one region to (spacelike, timelike) in other region, so that the names “time coordinate” and “radial coordinate” should be taken with a grain of salt.
@@
Regularity of the metric is defined in a sense that the associated Kretschmann scalar $K\mathrel{\mathop{:}}=R_{abcd}R^{abcd}$ is a bounded continuous real function on the interval $\left[0,\infty\right$>. In particular, we say that a metric has a “regular center” if $\lim_{r\to 0^{+}}K$ is well defined and finite. Kretschmann scalar $K$ for the metric (7) is a sum of three nonnegative terms,
@@
Note that a bounded Kretschmann scalar implies that left-hand sides of both gravitational field equations remain bounded as $r\to 0^{+}$. As was shown by two theorems in [BSJ22b], strictly stationary, regular solutions of the Einstein-NLE field equations, with the energy-momentum tensor respecting the dominant energy condition, have either a trivial or stealth electromagnetic field. Thus, our primary concern is with the regular black hole solutions.
@@ 4 Mass and charges entangled
In other words, to ensure that the Kretschmann scalar is bounded, the Komar mass $M$ and the magnetic charge $P$ must be constrained by the equation above. A special, pathological case occurs if $I=0$, implying that we need $M=0$ for regularization. Relation (22) should not come as a surprise, as it was already present in Ayón-Beato–García’s NLE interpretation of the Bardeen’s black hole (see comments in Appendix B) and it has appeared in more recent reverse engineered NLE Lagrangians [BFJS24] (see comments in Appendix A).
@@ 5 Generalized constraints
Let us turn to another kind of constraint on regular black holes with NLE fields. Back in 2001 Bronnikov [Bro01] proved that a NLE Lagrangian $\mathscr{L}(\mathcal{F})$ with the MWF limit does not admit electrically charged, static, spherically symmetric black holes with bounded Ricci scalar $R$ and “Ricci squared” scalar $S\mathrel{\mathop{:}}=R_{ab}R^{ab}$. This result was later generalized [BSJ22a] to a larger class of $\mathscr{L}(\mathcal{F},\mathcal{G})$ Lagrangians. Using the terminology from section 2, in combination with the recent results from [DMS25], it is straightforward to extend and strengthen the conclusion even further, in the form of the following theorem.
@@
This generalization immediately applies to theories respecting qMWF and not necessarily the MWF limit, such as ModMax theory. Another perspective may be added here, as ModMax theory belongs to the larger class of NLE theories with vanishing trace of energy-momentum tensor, $T=0$. Einstein’s gravitational field equation with the cosmological constant $\Lambda$ (included here for the sake of generality) implies for any member of such a theory that the Ricci scalar is constant, so that for static, spherically symmetric spacetimes considered here, we have